A mass weighting 32 lbs stretches a spring 8 inches. The mass is in a medium that exerts a viscous resistance of 24 lbs when the mass has a velocity of 2 ft/sec.

Suppose the object is displaced an additional 8 inches and released.

Find an equation for the object's displacement, u(t), in feet after t seconds.
Answer #1

# W = 32lbs., z=8 anch= 8 geht 32x12 lbs/ft = 48 lbs/ft 12 W = kx o 8 friscono = c 24 2 =12 Absos ft xlo= i (0)=0 .ft 12 ca &The complementary function m2 + 12.675m 48.3=0 me - 6.0375 +3 uur m2 = - 6.0375 -3.442 + Bemat mit . The soln = x(+) = 4e x=